OpenStudy (anonymous):

1. Two angles are complementary. The smaller angle is 35° less than the larger angle. What is the measure of the larger angle? A) 27.5° B) 62.5° C) 125° 2. A right circular cylinder has a base diameter of 12 meters and a height of 16 meters. What is the approximate total lateral surface area of the cylinder? A) 602.9 m² B) 1507.2 m² C) 2034.7 m²

OpenStudy (anonymous):

For number 1, the answer is B 62.5° but I don't know how you got that number:( and for number 2, does this mean that I have to add the lateral area and the surface area of the cylinder?

hartnn (hartnn):

complementary angles add upto 90 degrees so, if smaller angle is x, larger angle is x+35 right ?

hartnn (hartnn):

then just add them and equate to 90.

OpenStudy (anonymous):

@hartnn umm..ok I am totally blacked out..so, x+35 how would I solve for this..??

hartnn (hartnn):

smaller angle =x, larger angle =x+35 sum = 90 = x+x+35 so, can u find x from the above equation ?

OpenStudy (anonymous):

@hartnn oh~~!! yupp!! got it! it's 90=x+x+35 90=2x+35 55=2x 27.5=x so, the larger angle= x+35 27.5+35 = 62.5 right?!

hartnn (hartnn):

good work! correct :)

OpenStudy (anonymous):

@hartnn thank you!!! and if you can, can you help me out with the second question please?

hartnn (hartnn):

sure, only lateral surface area is asked. know the formula ?

OpenStudy (anonymous):

For the surface area, I know it's S.A.=2(PI)r*h+2(PI)r^2 and the lateral area...I'm not sure..

hartnn (hartnn):

total surface is made of 2 base areas, 2pi r^2 and lateral surface are 2pi r *h

OpenStudy (anonymous):

@hartnn is the "surface area" and the "total surface area" the same thing?

hartnn (hartnn):

yes.

hartnn (hartnn):

but theres the word lateral, so u need to use 2pi r*h

OpenStudy (anonymous):

@hartnn ah ok give me a sec, i'll solve for the surface area and the lateral area=)

OpenStudy (anonymous):

@hartnn ok so the Surface area is S.A.= 2(pi)r^2 S.A.= 2(pi)6^2 S.A.= 72 (pi) = 226.1946711 the lateral area is L.A.=2(pi)r*h L.A.=2(pi)6*16 L.A.=192(pi)=603.1857895

hartnn (hartnn):

LA was asked and its 603 is correct. Also , total surface area is always > lateral. what you calculated as 2pi r^2 was the total base area

hartnn (hartnn):

and SA =base area + lateral area.

hartnn (hartnn):

S.A.=2(PI)r*h+2(PI)r^2

OpenStudy (anonymous):

oh oh ok so, the surface area is 226.1946711+603.1857895 =829.3804606

hartnn (hartnn):

yes, correct.

hartnn (hartnn):

thats total surface area.

OpenStudy (anonymous):

ah so now should I add 829.3804606 and 603.1857895 together?

hartnn (hartnn):

why ?

hartnn (hartnn):

only lateral area is asked and its 603

OpenStudy (anonymous):

@hartnn oh I thought that (total "lateral" "surface" area of the cylinder) meant that I had to add the lateral area and surface area.

OpenStudy (anonymous):

@hartnn so they are just asking about the lateral area? then the answer would be A) 602.9 m² ?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

@hartnn Oh~ ok I see:) thank you soooo much!! you helped me a lot!!

hartnn (hartnn):

\[ \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{^}}\color{#e6e600}{\text{_}}\color{green}{\text{^}}\color{blue}{\text{}}\end{array} \]